OpenStudy (anonymous):

lim of (x+2)/(x^3+8) as x approaches -2

OpenStudy (anonymous):

\[=\lim_{x \rightarrow -2} \frac{x+2}{x^3 + 8}\]\[=\lim_{x \rightarrow -2}\frac{x+2}{(x+2)(x^2-4x+4)}\]\[=\lim_{x \rightarrow -2}\frac{1}{x^2-4x+4}\]\[=\frac{ 1}{ 4 + 8 + 4 }\]\[=\frac{ 1 }{ 12 }\]

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

wait, shouldn't it be equal to 1/16 because 4+8+4 = 16? where did you get 1/12 from?